OpenStudy (anonymous):

x^2+2x-15=0 solve by using the quadratic equation

OpenStudy (anonymous):

try factoring. it should work in this case

OpenStudy (anonymous):

could you show me how to use the quadratic on an equation like this, i know the answer by factoring but need help with the quadratic

OpenStudy (anonymous):

yeah here's the answer: quadratic equation:  plug in 2 as b; 1 as 4;15 as c. what you should get is \[-2\pm \sqrt{4+60}\]/ 2(1) \[(-2\pm \sqrt{64})/ 2 \] this becomes \[(-2\pm8)/2\] (-2+8)/ 2 = 3 and (-2-8)/2 = -5